OpenStudy (anonymous):

Which ratio forms a proportion with 15/21 ? 20/26, 7/5, 20/28, 12/8

OpenStudy (lgbasallote):

15/21 simplified is \[\frac{15 \div 3}{21 \div 3} \implies \frac 57\] so which one of those simplifies to 5/7?

OpenStudy (anonymous):

Precisely what @@lgbasallote said.

OpenStudy (anonymous):

20/28?

OpenStudy (lgbasallote):

\[\frac{20 \div 4}{28 \div 4} \implies \frac 57\\ \textbf{congrats!!}\]

OpenStudy (anonymous):

i mean 20/28

OpenStudy (anonymous):

thanks :) can you help me with another question?

OpenStudy (anonymous):

Which expression simplifies to 6h^2√3h? sqrt36h^5, sqrt49h^3, sqrt108h^5, sqrt18h^3